Write the point-slope form of the line's equation satisfying the given conditions. Then use the point-slope form of the equation to write the slope-intercept form. Slope m=-5. passing through (7,4)
Answer by rfer(16322)

y=mx+b
4=-5*7+b
4=-35+b
b=39
---------------
y=-5x+39